# The incompatibility of the Condorcet winner and loser criteria with positive or negative involvement and resolvability

### v4

@@ 3 Result
The case of positive involvement is given by Theorem 4. For negative involvement, first observe that a variant of Lemma 3 holds for negative involvement (cf. Pérez 2001, Lemma 1): Let $F$ be a voting method satisfying negative involvement and the Condorcet winner criterion. Let $\mathbf{P}$ be a profile in which for all $x,x^{\prime},y,y^{\prime}\in X(\mathbf{P})$ with $x\neq x^{\prime}$ and $y\neq y^{\prime}$, if $\text{Margin}_{\mathbf{P}}(x,x^{\prime})>\text{Margin}_{\mathbf{P}}(y,y^{\prime})$, then $\text{Margin}_{\mathbf{P}}(x,x^{\prime})>\text{Margin}_{\mathbf{P}}(y,y^{\prime})+1$. Where $j$ is the largest margin in $\mathcal{M}(\mathbf{P})$, additionally assume that $\mathbf{P}$ contains for each linear order of $X(\mathbf{P})$, $j+1$ voters submitting that linear order. [^7] This is a gratuitous assumption that can be weakened in an obvious way, but optimization is not necessary for the proof. Then $F(\mathbf{P})\subseteq D(\mathbf{P})$. The proof is the same as that of Lemma 3 except that instead of obtaining $\mathbf{P}^{\prime}$ from $\mathbf{P}$ by adding $k+1$ voters who rank $x$ uniquely first and $y$ uniquely second, followed by any linear order of $X(\mathbf{P})\setminus\{x,y\}$, we obtain $\mathbf{P}^{\prime}$ from $\mathbf{P}$ by removing $k+1$ voters who submit any linear order of $X(\mathbf{P})\setminus\{x,y\}$, followed by $y$ in the penultimate position, followed by $x$ uniquely last; these voters exist in $\mathbf{P}$ by our additional assumption on $\mathbf{P}$. Then since the proof of Lemma 3 begins with $x\in F(\mathbf{P})$, repeated application of negative involvement in its contrapositive form implies $x\in F(\mathbf{P}^{\prime})$. Removing a ranking $L$ has the same effect on margins as adding the reversed ranking $L^{-1}$, so as in the proof of Lemma 3, $y$ is the Condorcet winner in $\mathbf{P}^{\prime}$, yielding a contradiction.
Now we adapt the proof of Theorem 4 by adding 150 copies of each linear order of $X(\mathbf{P}_{1})$ to $\mathbf{P}_{1}$, which will ensure that the variant of Lemma 3 is applicable when needed in the proof. [^8] One can use an integer linear program to synthesize a new starting profile with far fewer voters such that all constraints needed for the proof (and for an optimized version of the variant of Lemma 3) still hold. Then instead of adding 26 voters with the ranking adbecadbec to obtain $\mathbf{P}_{2}$, we remove 26 voters with the ranking cebdacebda; instead of removing 7 voters with the ranking dacebdaceb from $\mathbf{P}_{2}$ to obtain $\mathbf{P}_{3}$, we add 7 voters with the ranking becadbecad; and so on for $\mathbf{P}_{4}$ (removing 23 voters with caedbcaedb) and $\mathbf{P}_{5}$ (adding 7 voters with ecabdecabd). The margins of $\mathcal{M}_{1},\dots,\mathcal{M}_{5}$ remain as in the proof of Theorem 4. Then the rest of the proof is the same as that of Theorem 4 except that we (i) invoke the variant of Lemma 3 for steps involving defensibility and (ii) use the contrapositive form of negative involvement in place of each application of positive involvement.∎
@@
For each positive integer $n$, there is no voting method satisfying the Condorcet winner and loser criteria, positive or negative involvement, and $n$-voter resolvability.
Thus, for any voting method satisfying the Condorcet winner and loser criteria and positive or negative involvement, no matter the size of $n$, sometimes a tied winner cannot be made a unique winner even by the addition of up to $n$ voters.
@@ Acknowledgements
I thank Emma Boniface and Diego Bejarano for helpful discussions of the problem posed at the end of Holliday 2024 and Dominik Peters for helpful feedback on this paper.

### v5

@@ 3 Result
A variant of Lemma 3 holds for negative involvement (cf. Pérez 2001, Lemma 1): Let $F$ be a voting method satisfying negative involvement and the Condorcet winner criterion. Let $\mathbf{P}$ be a profile in which for all $x,x^{\prime},y,y^{\prime}\in X(\mathbf{P})$ with $x\neq x^{\prime}$ and $y\neq y^{\prime}$, if $\text{Margin}_{\mathbf{P}}(x,x^{\prime})>\text{Margin}_{\mathbf{P}}(y,y^{\prime})$, then $\text{Margin}_{\mathbf{P}}(x,x^{\prime})>\text{Margin}_{\mathbf{P}}(y,y^{\prime})+1$. Additionally assume that for each $x\in X(\mathbf{P})\setminus D(\mathbf{P})$, there is a $y\in X(\mathbf{P})$ such that for every $z\in X(\mathbf{P})$, $\text{Margin}_{\mathbf{P}}(z,y)<\text{Margin}_{\mathbf{P}}(y,x)$, and where $k=\max\{\text{Margin}_{\mathbf{P}}(z,y)\mid z\in X(\mathbf{P}),z\neq y\}$, there are at least $k+1$ voters whose rankings have $y$ uniquely second-to-last, followed by $x$ uniquely last. Then $F(\mathbf{P})\subseteq D(\mathbf{P})$. The proof is the same as that of Lemma 3 except we obtain $\mathbf{P}^{\prime}$ from $\mathbf{P}$ by removing $k+1$ voters of the kind described above. Then since the proof of Lemma 3 begins with $x\in F(\mathbf{P})$, repeated application of negative involvement in its contrapositive form implies $x\in F(\mathbf{P}^{\prime})$. Removing a ranking $L$ has the same effect on margins as adding the reversed ranking $L^{-1}$, so as in the proof of Lemma 3, $y$ is the Condorcet winner in $\mathbf{P}^{\prime}$, yielding a contradiction.
Now we adapt the proof of Theorem 4 by adding 147 copies of each linear order of $X(\mathbf{P}_{1})$ to $\mathbf{P}_{1}$, which ensures that the variant of Lemma 3 is applicable when needed in the proof. [^6] By starting with at least $147$ copies of each linear order in $\mathbf{P}_{1}$, we ensure that there are at least $121$ copies of each linear order in $\mathbf{P}_{3}+\mathbf{S}_{2}$ and in $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$, which is greater than the maximal margins in those profiles, so the variant of Lemma 3 is applicable to those profiles. Of course, one can use an integer linear program to synthesize a new starting profile with far fewer voters such that all constraints needed for the proof still hold. Then instead of adding 26 voters with the ranking adbecadbec to obtain $\mathbf{P}_{2}$, we remove 26 voters with cebdacebda; instead of removing 7 voters with dacebdaceb from $\mathbf{P}_{2}$ to obtain $\mathbf{P}_{3}$, we add 7 voters with becadbecad; and so on for $\mathbf{P}_{4}$ (removing 23 voters with caedbcaedb) and $\mathbf{P}_{5}$ (adding 7 voters with ecabdecabd). The margins of $\mathcal{M}_{1},\dots,\mathcal{M}_{5}$ remain as in the proof of Theorem 4, and then the rest of the proof is the same except that we (i) invoke the variant of Lemma 3 for steps involving defensibility and (ii) use negative involvement instead of positive involvement.∎
@@
For each positive integer $n$, there is no voting method satisfying the Condorcet winner and loser criteria, negative involvement, and $n$-voter resolvability.
Finally, in response to a previous version of this note with the results above, Dominik Peters (personal communication) suggested that analogous impossibilities might hold with independence of clones (Tideman 1987) instead of the Condorcet loser criterion. Indeed, we can confirm that such a swap in Theorems 4 and 6 is possible, as well as in Theorems 78 if we add the axiom of block preservation (Holliday and Pacuit 2025): if the profile $\mathbf{L}$ consists of exactly one copy of each linear order of $X(\mathbf{P})$, then $F(\mathbf{P})\subseteq F(\mathbf{P}+\mathbf{L})$.
Recall that a set $C\subsetneq X(\mathbf{P})$ with $|C|\geq 2$ is a set of clones in $\mathbf{P}$ if for each $i\in V(\mathbf{P})$ and $x\in X(\mathbf{P})\setminus C$, either $i$ ranks $x$ above each $c\in C$ or $i$ ranks $x$ below each $c\in C$. Given a profile $\mathbf{P}$ and $c\in X(\mathbf{P})$, let $\mathbf{P}_{-c}$ be the profile with $X(\mathbf{P}_{-c})=X(\mathbf{P})\setminus\{c\}$ and $V(\mathbf{P}_{-c})=V(\mathbf{P})$ where for each $i\in V(\mathbf{P})$, $\mathbf{P}_{-c}(i)$ is the restriction of $\mathbf{P}(i)$ to $X(\mathbf{P}_{-c})$. Then $F$ satisfies independence of clones if for any profile $\mathbf{P}$, set $C$ of clones, $c\in C$, and $x\in X(\mathbf{P})\setminus C$, (i) $x\in F(\mathbf{P})$ iff $x\in F(\mathbf{P}_{-c})$, and (ii) $C\cap F(\mathbf{P})\neq\varnothing$ iff $C\cap F(\mathbf{P}_{-c})\neq\varnothing$.
Theorem 9.
For each positive integer $n$, there is no voting method satisfying the Condorcet winner criterion, independence of clones, positive involvement, and $n$-voter resolvability.
The same holds with positive involvement replaced by negative involvement plus block preservation.
Proof.
We prove the $n=1$ case. Reasoning as in the proof of Theorem 6 handles $n>1$. For positive involvement, consider a $\mathbf{Q}_{1}$ as in Figure 3, wherein $\{a,b,c,e\}$ is a set of clones. By Lemma 3, $F(\mathbf{Q}_{1})\subseteq\{a,d\}$. By the Condorcet winner criterion, $F(\mathbf{Q}_{1-b})=\{a\}$, so by independence of clones (part (i) or (ii)) and $F(\mathbf{Q}_{1})\subseteq\{a,d\}$, we have $a\in F(\mathbf{Q}_{1})$. Then obtain $\mathbf{Q}_{2},\dots,\mathbf{Q}_{5}$ from $\mathbf{Q}_{1}$ by adding and subtracting the profiles in Figure 4. The rest of the proof is just like that of Theorem 4 but with $\mathbf{Q}_{i}$ instead of $\mathbf{P}_{i}$. For negative involvement, by block preservation, we can add as many blocks of all linear orders as we wish to $\mathbf{Q}_{1}$ to obtain a $\mathbf{Q}_{1}^{+}$ with $F(\mathbf{Q}_{1})\subseteq F(\mathbf{Q}_{1}^{+})\subseteq\{a,d\}$ by the variant of Lemma 3 in the proof of Theorem 7. Hence $F(\mathbf{Q}_{1})\subseteq\{a,d\}$, which implies $a\in F(\mathbf{Q}_{1})$ as above and therefore $a\in F(\mathbf{Q}_{1}^{+})$. Then starting from $\mathbf{Q}_{1}^{+}$, we reason with negative involvement instead of positive involvement as in the proof of Theorem 7.∎
62 60 42 23 19 3
$d$ $c$ $d$ $a$ $e$ $c$
$b$ $b$ $e$ $e$ $c$ $e$
$a$ $a$ $a$ $c$ $b$ $b$
$e$ $e$ $c$ $b$ $a$ $a$
$c$ $d$ $b$ $d$ $d$ $d$
Figure 3: the profile $\mathbf{Q}_{1}$ used in the proof of Theorem 9.
$\mathcal{M}_{1}$
aabbccddee45451181817979113535858511838311
+\;+ 16 7 2 $a$ $a$ $a$ $d$ $e$ $e$ $b$ $d$ $b$ $c$ $b$ $d$ $e$ $c$ $c$ ==\;\; $\mathcal{M}_{2}$aabbccddee7070262610610654544242606020202424667676 == 5 $d$ $e$ $a$ $c$ $b$ ++\;\; $\mathcal{M}_{3}$aabbccddee6565313111111159594747555515151919117171
$\mathcal{M}_{3}$
aabbccddee6565313111111159594747555515151919117171
+\,\;\;\;\,+\;\;\; 20 $b$ $d$ $c$ $a$ $e$ =\;\;\;=\;\;\;\;\;\,\, $\mathcal{M}_{4}$aabbccddee4545111113113179795567673535393921215151 == 11 $d$ $b$ $a$ $e$ $c$ ++\;\; $\mathcal{M}_{5}$aabbccddee343422221201206868161656564646282810104040
Figure 4: margin graphs used in the proof of Theorem 9 with defensible candidates shaded gray.
@@ Acknowledgements
I thank Diego Bejarano and Emma Boniface for helpful discussions of the problem posed at the end of Holliday 2024 and Dominik Peters for helpful feedback on this paper.